Given A= i+2j+3k and B= 2i+j-k, find AxB


1
Expert's answer
2021-06-30T16:53:42-0400
"A\times B=\begin{vmatrix}
 i & j & k \\
 1 & 2 & 3 \\
2 & 1 & -1
\end{vmatrix}"

"=i\begin{vmatrix}
 2 & 3 \\
 1 & -1
\end{vmatrix}-j\begin{vmatrix}
 1 & 3 \\
 2 & -1
\end{vmatrix}+k\begin{vmatrix}
 1 & 2 \\
 2 & 1
\end{vmatrix}"

"=i(-2-3)-j(-1-6)+k(1-4)"

"=-5i+7j-3k"


"A\times B=-5i+7j-3k"